THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 101 — 5 October 2009

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues, It is my pleasure to present you the 101^{st} issue of the Magellanic Clouds Newsletter. Interesting new work is presented on the star formation and molecular clouds in the Magellanic Clouds, X-ray observations, and star clusters, and more. Good news for people on the hunt for a job: a postdoc position has opened in Brussels, to work on the VISTA near-infrared survey of the Magellanic Clouds (VMC), which is about to commence. The next issue is planned to be distributed on the 1^{st} of December 2009.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

High and intermediate-mass young stellar objects in the Large Magellanic Cloud

Robert A. Gründl^{1} and You-Hua Chu^{1,2}

1University of Illinois, USA 2Visiting astronomer, Cerro Tololo Inter-American Observatory

Archival Spitzer IRAC and MIPS observations of the Large Magellanic Cloud (LMC) have been used to search for young stellar objects (YSOs). We have carried out independent aperture photometry of these data and merged the results from different passbands to produce a photometric catalog. To verify our methodology we have also analyzed the data from the SAGE and SWIRE Legacy programs; our photometric measurements are in general agreement with the photometry released by these programs. A detailed completeness analysis for our photometric catalog of the LMC show that the 90% completeness limits are, on average, 16.0, 15.0, 14.3, 13.1, and 9.2 mag at 3.6, 4.5, 5.8, 8.0, and 24 μm, respectively. Using our mid-infrared photometric catalogs and two simple selection criteria, [4.5]−[8.0] > 2.0 to exclude normal and evolved stars and [8.0] > 14−([4.5]−[8.0]) to exclude background galaxies, we have identified a sample of 2,910 sources in the LMC that could potentially be YSOs. We then used the Spitzer observations complemented by optical and nearinfrared data to carefully assess the nature of each source. To do so we simultaneously considered multi-wavelength images and photometry to assess the source morphology, spectral energy distribution (SED) from the optical through the mid-infrared wavelengths, and the surrounding interstellar environment to determine the most likely nature of each source. From this examination of the initial sample, we suggest 1,172 sources are most likely YSOs. We have also identified 1,075 probable background galaxies, consistent with the expected number estimated from the SWIRE survey. Spitzer IRS observations of 269 of the brightest YSOs from our sample have confirmed that > ∼95% are indeed YSOs. Examination of color–color and color–magnitude diagrams shows no simple criteria in color–magnitude space that can unambiguously separate the LMC YSOs from all AGB/post-AGB stars, planetary nebulae, and background galaxies. A comprehensive search for YSOs in the LMC has also been carried out by the SAGE team and reported by Whitney et al. (2008). There are three major differences between these two searches. (1) In the common region of color–magnitude space, ∼ 850 of our 1,172 probable YSOs are missed in the SAGE YSO catalog because their conservative point source identification criteria have excluded YSOs superposed on complex stellar and interstellar environments. (2) About 20–30% of the YSOs identified by the SAGE team are sources we classify as background galaxies. (3) the SAGE YSO catalog identifies YSO in parts of color–magnitude space that we excluded and thus contains more evolved or fainter YSOs missed by our analysis. The shortcomings and strengths of both these YSO catalogs should be considered prior to statistical studies of star formation in the LMC. Finally, the mid-infrared luminosity functions in the IRAC bands of our most likely YSO candidates in the LMC can be well described by N(L) ∝ L^{−1}, which is consistent with the Salpeter initial mass function if a mass-luminosity relation of L ∝ M^{2.4} is adopted.

Accepted for publication in Astrophysical Journal Supplement Series Available from arXiv:0908.0347 and from http://www.astro.illinois.edu/∼gruendl/LMC YSO.pdf

The star formation history of the Large Magellanic Cloud

Jason Harris^{1} and Dennis Zaritsky^{2}

1NOAO, USA 2University of Arizona, USA

We present the first-ever global, spatially-resolved reconstruction of the star formation history (SFH) of the Large Magellanic Cloud (LMC), based on the application of our StarFISH analysis software to the multiband photometry of twenty million of its stars from the Magellanic Clouds Photometric Survey. The general outlines of our results are consistent with previously published results: following an initial burst of star formation, there was a quiescent epoch from ≈ 12–5 Gyr ago. Star formation then resumed and has proceeded until the current time at an average rate of roughly 0.2 M_{⊙} yr^{−1}, with temporal variations at the factor-of-two level. The re-ignition of star formation ≈ 5 Gyr ago, in both the LMC and SMC, is suggestive of a dramatic event at that time in the Magellanic system. Among the global variations in the recent star formation rate are peaks at roughly 2 Gyr, 500 Myr, 100 Myr and 12 Myr. The peaks at 500 Myr and 2 Gyr are nearly coincident with similar peaks in the SFH of the Small Magellanic Cloud, suggesting a joint history for these galaxies extending back at least several Gyr. The chemical enrichment history recovered from our StarFISH analysis is in broad agreement with that inferred from the LMC’s star cluster population, although our constraints on the ancient chemical enrichment history are weak. We conclude from the concordance between the star formation and chemical enrichment histories of the field and cluster populations that the field and cluster star formation modes are tightly coupled.

Accepted for publication in AJ Available from arXiv:0908.1422

Infrared dark clouds in the Small Magellanic Cloud?

Min-Young Lee^{1}, Snežana Stanimirović^{1}, Jürgen Ott^{2}, Jacco Th. van Loon^{3}, Alberto D. Bolatto^{4}, Paul A. Jones^{5,6}, Maria R. Cunningham^{5}, Kathryn E. Devine^{1} and Joana M. Oliveira^{4}

1Department of Astronomy, University of Wisconsin at Madison, USA 2National Radio Astronomy Observatory, USA 3Astrophysics Group, Lennard Jones Laboratories, Keele University, UK 4Department of Astronomy and Laboratory for Millimeter-wave Astronomy, University of Maryland, College Park, USA 5School of Physics, University of New South Wales, Australia 6Departamento de Astronomía, Universidad de Chile, Chile

We have applied the unsharp-masking technique to the 24 μm image of the Small Magellanic Cloud (SMC), obtained with the Spitzer Space Telescope, to search for high-extinction regions. This technique has been used to locate very dense and cold interstellar clouds in the Galaxy, particularly infrared dark clouds (IRDCs). Fifty-five candidate regions of high extinction, namely high-contrast regions (HCRs), have been identified from the generated decremental contrast image of the SMC. Most HCRs are located in the southern bar region and mainly distributed in the outskirts of CO clouds, but most likely contain a significant amount of H_{2}. HCRs have a peak-contrast at 24 μm of 2–2.5% and a size of 8–14 pc. This corresponds to the size of typical and large Galactic IRDCs, but Galactic IRDCs are 2–3 times darker at 24 μm than our HCRs. To constrain the physical properties of the HCRs, we have performed NH_{3}, N_{2}H^{+}, HNC, HCO^{+}, and HCN observations toward one of the HCRs, HCR LIRS36-east, using the Australia Telescope Compact Array and the Mopra single-dish radio telescope. We did not detect any molecular line emission, however, our upper limits to the column densities of molecular species suggest that HCRs are most likely moderately dense with n ∼ 10^{3} cm^{−3}. This volume density is in agreement with predictions for the cool atomic phase in low metallicity environments. We suggest that HCRs may be tracing clouds at the transition from atomic to molecule-dominated medium, and could be a powerful way to study early stages of gas condensation in low metallicity galaxies. Alternatively, if made up of dense molecular clumps < 0.5 pc in size, HCRs could be counterparts of Galactic IRDCs, and/or regions with highly unusual abundance of very small dust grains.

Accepted for publication in The Astronomical Journal Available from arXiv:0908.2275

XMMSL1 J060636.2−694933: An XMM-Newton Slew discovery and Swift/Magellan follow up of a new Classical Nova in the LMC

A.M. Read^{1}, R.D. Saxton^{2}, P.G. Jonker^{3,4}, E. Kuulkers^{2}, P. Esquej^{1}, G. Pojmański^{5}, M.A.P. Torres^{4}, M.R. Goad^{1}, M.J. Freyberg^{6} and M. Modjaz^{7}

1Dept. of Physics and Astronomy, Leicester University, Leicester LE1 7RH, U.K. 2XMM-Newton SOC, ESAC, Apartado 78, 28691 Villanueva de la Cañada, Madrid, Spain 3SRON, Netherlands Institute for Space Research, Sorbonnelaan 2, 3584 CA, Utrecht, The Netherlands 4Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138, U.S.A. 5Warsaw University Observatory, A1. Ujazdowskie 4, 00-478, Warsaw, Poland 6Max-Planck-Institut fuer extraterrestrische Physik, 85748 Garching, Germany 7University of California, 601 Campbell Hall, Berkeley, CA 94720

In order to discover new X-ray transients, the data taken by XMM-Newton as it slews between targets are being processed and cross-correlated with other X-ray observations. A bright source, XMMSL1 J060636.2−694933, was detected on 18 July 2006 at a position where no previous X-ray source had been seen. The XMM-Newton slew data, plus follow-up dedicated XMM-Newton and Swift observations, plus optical data acquired with the Magellan Clay telescope, and archival All-Sky Automated Survey (ASAS) data were used to classify the new object, and to investigate its properties. No XMM-Newton slew X-ray counts are detected above 1 keV and the source is seen to be over five hundred times brighter than the ROSAT All-Sky Survey upper limit at that position. The line-rich optical spectrum acquired with the Magellan telescope allows the object to be classified as an Ao auroral phase nova, and the soft X-ray spectrum indicates that the nova was in a super-soft source state in the X-ray decline seen in the follow-up X-ray observations. The archival ASAS data suggests that the nova at onset (Oct 2005) was a ’very fast’ nova, and an estimate of its distance is consistent with the nova being situated within the LMC. With the discovery presented here of a new classical nova in the LMC, it is clear that XMM-Newton slew data are continuing to offer a powerful opportunity to find new X-ray transient objects.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:0908.3989

LMC self lensing for OGLE-II microlensing observations

S. Calchi Novati^{1,2,3}, L. Mancini^{1,2,3}, G. Scarpetta^{1,2,3} and Ł. Wyrzykowski^{4}

1Dipartimento di Fisica, Università di Salerno, 84081 Baronissi (SA), Italy 2INFN, Sez. di Napoli, Italy 3IIASS, Italy 4Institute of Astronomy, University of Cambridge, Cambridge CB3 0HA, UK

In the framework of microlensing searches towards the Large Magellanic Cloud (LMC), we discuss the results presented by the OGLE collaboration for their OGLE-II campaign (Wyrzykowski et al. 2009). We evaluate the optical depth, the duration and the expected rate of events for the different possible lens populations: both luminous, dominated by the LMC self lensing, and “dark”, the would be compact halo objects (MACHOs) belonging to either the Galactic or to the LMC halo. The OGLE-II observational results, 2 microlensing candidate events located in the LMC bar region with duration of 24.2 and 57.2 days, compare well with the expected signal from the luminous lens populations: n_{exp} = 1.5, with typical duration, for LMC self lensing, of about 50 days. Because of the small statistics at disposal, however, the conclusions that can be drawn as for the halo mass fraction, f, in form of compact halo objects are not too severe. By means of a likelihood analysis we find an upper limit for f, at 95% confidence level, of about 15% in the mass range (10^{−2}–10^{−1}) M_{⊙} and 26% for 0.5 M_{⊙}.

Accepted for publication in MNRAS Available from arXiv:0908.3836

Multifrequency Observations of One of the Largest Supernova Remnants in the Local Group of Galaxies, LMC — SNR J0450−709

K.O. Čajko^{1}, E.J. Crawford^{2} and M.D. Filipović^{2}

1Faculty of Sciences, University of Novi Sad, Serbia 2University of Western Sydney, Australia

We present the results of new Australia Telescope Compact Array (ATCA) observations of one of the largest supernova remnants, SNR J0450−709, in the Local Group of galaxies. We found that this Large Magellanic Cloud (LMC) object exhibits a typical morphology of an old supernova remnant (SNR) with diameter D = 102 × 75 ± 1 pc and radio spectral index α = −0.43 ± 0.06. Regions of high polarisation were detected with peak value of ∼ 40%.

Accepted for publication in Serbian Astronomical Jounal Available from arXiv:0909.0310

Exploring accretion theory with X-ray binaries in the SMC

M.J. Coe^{1}, V.A. McBride^{1} and R.H.D. Corbet^{2}

1University of Southampton, UK 2NASA/GSFC, USA

The understanding of the accretion process on to compact objects in binary systems is an important part of modern astrophysics. Theoretical work, primarily that of Ghosh & Lamb (1979), has made clear predictions for the behaviour of such systems which have been generally supported by observational results of considerably varying quality from galactic accreting pulsar systems. In this work a much larger homogeneous population of such objects in the Small Magellanic Cloud (SMC) is used to provide more demanding tests of the accretion theory. The results are extremely supportive of the theoretical models and provide useful statistical insights into the manner in which accreting pulsars behave and evolve.

Accepted for publication in MNRAS Available from arXiv:0909.4332

Structural parameters of seven SMC intermediate-age and old star clusters

Katharina Glatt^{1,2}, Eva K. Grebel^{1,2}, John S. Gallagher^{3}, Antonella Nota^{4}, Elena Sabbi^{4}, Marco Sirianni^{4}, Gisella Clementini^{5}, Gary Da Costa^{6}, Monica Tosi^{5}, Daniel Harbeck^{3}, Andreas Koch^{7} and Andrea Kayser^{1}

1Department of Physics, University of Basel, Klingelbergstraße 82, CH-4056 Basel, Switzerland 2Astronomisches Rechen-Institut, Zentrum für Astronomie der Universität Heidelberg, Mönchhofstraße 12-14, D-69120 Heidelberg, Gernamy 3Department of Astronomy, University of Wisconsin, 475 North Charter Street, Madison, WI 53706-1582, USA 4Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 5INAF-Osservatorio Astronomica di Bologna, Via Ranzani 1, 40127 Bologna, Italy 6Research School of Astronomy & Astrophysics, The Australian National University, Mt. Stromlo Observatory, via Cotter Rd, Weston, ACT 2611, Australia 7Department of Physics and Astronomy, University of Leicester, University Road, Leicester, LE1 7RH, UK

We present structural parameters for the seven intermediate-age and old star clusters NGC 121, Lindsay 1, Kron 3, NGC 339, NGC 416, Lindsay 38, and NGC 419 in the Small Magellanic Cloud. We fit King profiles and Elson, Fall, & Freeman (EFF) profiles to both surface-brightness and star count data taken with the Advanced Camera for Surveys aboard the Hubble Space Telescope. Clusters older than ∼ 1 Gyr show a spread in cluster core radii that increases with age, while the youngest clusters have relatively compact cores. No evidence for post core collapse clusters was found. We find no correlation between core radius and distance from the SMC center, although consistent with other studies of dwarf galaxies, some relatively old and massive clusters have low densities. Clusters younger than 5 Gyr have larger ellipticities than older clusters, but the oldest SMC star cluster, the only globular NGC 121, is the most elliptical in our study. No correlation is seen between ellipticity and distance from the SMC center. The structures of these massive intermediate-age (1–8 Gyr) SMC star clusters thus appear to primarily result from internal evolutionary processes.

Accepted for publication in The Astronomical Journal Available from arXiv:0909.1329

A slitless spectroscopic survey for Hα emission-line objects in SMC clusters

Christophe Martayan^{1,2}, Dietrich Baade^{3} and Juan Fabregat^{4}

1European Organisation for Astronomical Research in the Southern Hemisphere, Alonso de Cordova 3107, Vitacura, Casilla 19001, Santiago 19, Chile 2GEPI, Observatoire de Paris, CNRS, Université Paris Diderot, 5 place Jules Janssen, 92195 Meudon Cedex, France 3European Organisation for Astronomical Research in the Southern Hemisphere, Karl-Schwarzschild-Str. 2, 85748 Garching b. München, Germany 4Observatorio Astronómico de Valencia, edifici Instituts d’investigació, Poligon la Coma, 46980 Paterna Valencia, Spain

Context: A fair fraction of all single early-type stars display emission lines well before the supergiant phase. Very rapid rotation is necessary for such stars to form rotationally supported decretion disks. But it is unknown whether and which other parameters may be important. Aims: This paper checks on the roles of metallicity and evolutionary age in the appearance of the so-called Be phenomenon. Methods: Slitless CCD spectra were obtained covering the bulk (about 3 square degrees) of the Small Magellanic Cloud. For Hα line emission twice as strong as the ambient continuum, the survey is complete to spectral type B2/B3 on the main sequence. About 8,120 spectra of 4,437 stars were searched for emission lines in 84 open clusters. 370 emission-line stars were found, among them at least 231 near the main sequence. For 176 of them, photometry could be found in the OGLE database. For comparison with a higher-metallicity environment, the Galactic sample of the photometric Hα survey by McSwain & Gies (2005) was used. Results: Among early spectral sub-types, Be stars are more frequent by a factor ∼3–5 in the SMC than in the Galaxy. The distribution with spectral type is similar in both galaxies, i.e. not strongly dependent on metallicity. The fraction of Be stars does not seem to vary with local star density. The Be phenomenon mainly sets in towards the end of the main-sequence evolution (this trend may be more pronounced in the SMC); but some Be stars already form with Be-star characteristics. In small sub-samples (such as single clusters), even if they appear identical, the fraction of emission lines stars can deviate drastically from the mean. Conclusions: In all probability, the fractional critical angular rotation rate, Ω/Ω_{c}, is one of the main parameters governing the occurrence of the Be phenomenon. If the Be character is only acquired during the course of evolution, the key circumstance is the evolution of Ω/Ω_{c}, which not only is dependent on metallicity but differently so for different mass ranges. As the result, even if the Be phenomenon is basically single-parametric (namely Ω/Ω_{c}), it takes on a complex multi-parametric appearance. The large cluster-to-cluster differences, which seem stronger than all other variations, serve as a caveat that this big picture may undergo significant second-order modulations (pulsations, initial angular momentum, etc.).

Accepted for publication in A&A Available from arXiv:0909.2303 and from http://hal.archives-ouvertes.fr/hal-00416129/fr/

The mid-infrared colours of Galactic Bulge, Disk and Magellanic Planetary Nebulae

J.P. Phillips^{1} and G. Ramos-Larios^{1}

1Instituto de Astronomía y Meteorología, Av. Vallarta No. 2602, Col. Arcos Vallarta, C.P. 44130 Guadalajara, Jalisco, México

We present mid-infrared (MIR) photometry for 367 Galactic disk, bulge and Large Magellanic Cloud (LMC) planetary nebulae, determined using GLIMPSE II and SAGE data acquired using the Spitzer Space Telescope. This has permitted us to make a comparison between the luminosity functions of bulge and LMC planetary nebulae, and between the MIR colours of all three categories of source. It is determined that whilst the 3.6 μm luminosity function of the LMC and bulge sources are likely to be closely similar, the [3.6]–[5.8] and [5.8]–[8-0] indices of LMC nebulae are different from those of their disk and bulge counterparts. This may arise because of enhanced 6.2 μm PAH emission within the LMC sources, and/or as a result of differences between the spectra of LMC PNe and those of their Galactic counterparts. We also determine that the more evolved disk sources listed in the MASH catalogues of Parker et al. and Miszalski et al. (2008) have similar colours to those of the less evolved (and higher surface brightness) sources in the catalogue of Acker et al. (1992); a result which appears at variance with previous studies of these sources. Published in MNRAS Available from arXiv:0909.2069

Old main-sequence turnoff photometry in the Small Magellanic Cloud. II. Star formation history and its spatial gradients

Noelia E. D. Noël^{1}, Antonio Aparicio^{2}, Carme Gallart^{2}, Sebastián L. Hidalgo^{2}, Edgardo Costa^{3} and René A. Méndez^{3}

1Institute for Astronomy, Royal Observatory, University of Edinburgh, UK 2Instituto de Astrofísica de Canarias, Spain 3Departamento de Astronomía, Universidad de Chile, Chile

We present a quantitative analysis of the star formation history (SFH) of 12 fields in the Small Magellanic Cloud (SMC) based on unprecedented deep [(B–R),R] color–magnitude diagrams (CMDs). Our fields reach down to the oldest main sequence (MS) turnoff with high photometric accuracy, which is vital for obtaining accurate SFHs, particularly at intermediate and old ages. We find that there are four main periods of enhancement of star formation: a young one peaked at around 0.2–0.5 Gyr old, only present in the eastern and in the central-most fields; two at intermediate ages present in all fields (a conspicuous one peaked at 4–5 Gyr old, and a less significant one peaked at 1.5–2.5); and an old one, peaked at 10 Gyr in all fields but the western ones. In the western fields, this old enhancement splits into two, one peaked at around 8 Gyr old and another at around 12 Gyr old. This “two-enhancement” zone seems to be a robust feature since it is unaffected by our choice of stellar evolutionary library but more data covering other fields of the SMC are necessary in order to ascertain its significancy. Correlation between the star formation rate enhancements and SMC–Milky Way encounters is not clear. Some correlation could exist with encounters taken from the orbit determination of Kallivayalil et al. (2006). But our results would be also fit in a first pericenter passage scenario like the one claimed by Besla et al. (2007). For SMC–Large Magellanic Cloud (LMC) encounters, we find a correlation only for the most recent encounter about 0.2 Gyr ago. The population younger than 1 Gyr old in the wing area represents about 7–12% of the total SFR. This is significant in the sense that these eastern fields are the only ones of this study in which star formation is currently going on. There is a strong dichotomy between East/Southeast and West in the current irregular shape of the SMC. We find that this dichotomy is produced by the youngest population and began about 1 Gyr ago or later. The age of the old population is similar at all radii and at all azimuth and we constrain the age of this oldest population to be older than around 12 Gyr old. We do not find yet a region dominated by an old, Milky Way-like, halo at 4.5 kpc from the SMC center, indicating either that this old stellar halo does not exist in the SMC or that its contribution to the stellar populations, at the galactocentric distances of our outermost field, is negligible. Finally, we derive the age–metallicity relation and find that, in all fields, the metallicity increased continuously from early epochs until the present. Accepted for publication in Astronomical Journal Available from arXiv:0909.2873 and from http://www.iac.es/galeria/noelia/PaperII Figures/index.html (high-resolution paper and figures)

The influence of far-ultraviolet radiation on the properties of molecular clouds in the 30 Dor region of the Large Magellanic Cloud

Jorge L. Pineda^{1}, Jürgen Ott^{2}, Tony Wong^{3}, Erik Muller^{4} and Annie Hughes^{5}

1Argelander Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2National Radio Astronomy Observatory, P.O. Box O, Socorro, NM 87801, USA 3School of Physics, University of New South Wales, Sydney NSW 2052, Australia 4Department of Astrophysics, Nagoya University, Furo-cho, Chikusa-ku, Nagoya 464-8602, Japan 5Centre for Supercomputing and Astrophysics, Swinburne University of Technology, Hawthorn, VIC 3122, Australia

We present a complete ^{12}CO J = 1 → 0 map of the prominent molecular ridge in the Large Magellanic Cloud (LMC) obtained with the 22-m ATNF Mopra Telescope. The region stretches southward by ∼ 2^{∘} (or 1.7 kpc) from 30 Doradus, the most vigorous star-forming region in the Local Group. The location of this molecular ridge is unique insofar as it allows us to study the properties of molecular gas as a function of the ambient radiation field in a low-metallicity environment. We find that the physical properties of CO-emitting clumps within the molecular ridge do not vary with the strength of the far-ultraviolet (FUV) radiation field. Since the peak CO brightness of the clumps shows no correlation with the radiation field strength, the observed constant value for CO-to-H_{2} conversion factor along the ridge seems to require an increase in the kinetic temperature of the molecular gas that is offset by a decrease in the angular filling factor of the CO emission. We find that the difference between the CO-to-H_{2} conversion factor in the molecular ridge and the outer Milky Way is smaller than has been reported by previous studies of the CO emission: applying the same cloud identification and analysis methods to our CO observations of the LMC molecular ridge and CO data from the outer Galaxy survey by Dame et al. (2001), we find that the average CO-to-H_{2} conversion factor in the molecular ridge is X_{CO} ≃ (3.9 ± 2.5) × 10^{20} cm^{−2} (K km s^{−1})^{−1}, approximately twice the value that we determine for the outer Galaxy clouds. The mass spectrum and the scaling relations between the properties of the CO clumps in the molecular ridge are similar, but not identical, to those that have been established for Galactic molecular clouds.

Published in ApJ Available from arXiv:0907.5186

Conference Papers

IRSF/SIRIUS JHK_{s} near-infrared variable star survey in the Magellanic Clouds

Yoshifusa Ita^{1} and IRSF/SIRIUS variable star survey team

1National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo, 181-8588 Japan

We carried out a NIR variable star survey toward the Large and Small Magellanic Clouds using the InfraRed Survey Facility (IRSF) at Sutherland, South African Astronomical Observatory. This survey project was initiated in December 2000, and since then, we kept monitoring a total area of 3 square degrees along the LMC bar and also an area of 1 square degree around the center of the SMC, sufficiently large to do statistical analysis and to make complete catalog of variable red giants in the Magellanic Clouds. The detection limits (S/N = 10) of the survey are 17.0, 16.5 and 15.5 at J, H and K_{s}, respectively. In this article, we present some results on infrared variables that are not detected by the previous optical surveys. We show that they do not fall on the standard period–K_{s} magnitude relation for Mira-type variables pulsating in the fundamental mode.

Poster contribution, published in ”Stellar Pulsation: Challenges for Theory and Observation”, Eds. J. Guzik & P. Bradley Available from arXiv:0907.5244

Dense star-forming gas and dust in the Magellanic Clouds

Frank P. Israel^{1}

1Sterrewacht, Leiden University, Leiden, The Netherlands

The early stages of the star formation process are closely related to the condition of the parent interstellar medium, in particular to the heating/cooling balance, which itself is a function of ambient conditions. Important questions such as dust abundance, size distribution, temperature distribution, fraction of molecular gas, fraction of dense gas, gas surface density and total amount of gas and dust require separation of metallicity and radiation effects. The Magellanic Clouds provide an ideal laboratory to carry out such studies almost under ’controlled conditions’. Although they are prominent targets for space observatories (Spitzer, Herschel), an important role remains for large groundbased facilities, such as a 25-m class sub-millimeter telescope on Dome C. Large-scale mapping at high resolution should be carried out both in the continuum and in various lines, fully complementing other groundbased and spacebased observing programs. Oral contribution, published in 3^{rd} ARENA Conference: ”An Astronomical Observatory at Concordia”, Frascati (Italy), May 11–15 (2009) Available from arXiv:0908.3584

Multiphase PC/PL Relations: Comparison between Theory and observations

S. Kanbur^{1}, M. Marconi^{2}, C. Ngeow^{3}, I. Musella^{2}, M. Turner^{4}, S. Magin^{1}, J. Hasley^{1} and C. Bisse^{1}

1SUNY Oswego, USA 2Naples Observatory, Italy 3University of Illinois, USA 4Rice University, USA

Cepheids are fundamental objects astrophysically in that they hold the key to a CMB independent estimate of Hubble’s constant. A number of researchers have pointed out the possibilities of breaking degeneracies between Ω_{Matter} and H_{0} if there is a CMB independent distance scale accurate to a few percent (Hu 2005). Current uncertainties in the distance scale are about 10% but future observations, with, for example, the JWST, will be capable of estimating H_{0} to within a few percent. A crucial step in this process is the Cepheid PL relation. Recent evidence has emerged that the PL relation, at least in optical bands, is nonlinear and that neglect of such a nonlinearity can lead to errors in estimating H_{0} of up to 2 percent. Hence it is important to critically examine this possible nonlinearity both observationally and theoretically. Existing PC/PL relations rely exclusively on evaluating these relations at mean light. However, since such relations are the average of relations at different phases. Here we report on recent attempts to compare theory and observation in the multiphase PC/PL planes. We construct state of the art Cepheid pulsations models appropriate for the LMC/Galaxy and compare the resulting PC/PL relations as a function of phase with observations. For the LMC, the (V − I) period–color relation at minimum light can have quite a narrow dispersion (0.2–0.3 mags) and thus could be useful in placing constraints on models. At longer periods, the models predict significantly redder (by about 0.2–0.3 mags) V − I colors. We discuss possible reasons for this and also compare PL relations at various phases of pulsation and find clear evidence in both theory and observations for a nonlinear PL relation. Oral contribution, published in ”Stellar Pulsation: Challenges for Theory and Observation” Available from arXiv:0908.3947

The connection between pulsation, mass loss and circumstellar shells in

Classical Cepheids

Hilding Neilson^{1}, Chow-Choong Ngeow^{2}, Shashi Kanbur^{3} and John B. Lester^{4}

1University of Toronto, Canada 2UIUC, USA 3SUNY Oswego, USA 4University of Toronto Mississauga, Canada

Recent observations of Cepheids using infrared interferometry and Spitzer photometry have detected the presence of circumstellar envelopes (CSE) of dust and it has been hypothesized that the CSEs are due to dust forming in a Cepheid wind. Here we use a modified Castor, Abbott & Klein formalism to produce a Cepheid wind, and this is used to estimate the contribution of mass loss to the Cepheid mass discrepancy Furthermore, we test the OGLE-III Classical Cepheids using the IR fluxes from the SAGE survey to determine if Large Magellanic Cloud Cepheids have CSE’s. It is found that IR excess is a common phenomenon for LMC Cepheids and that the resulting mass-loss rates can explain at least a fraction of the Cepheid mass discrepancy, depending on the assumed dust-to-gas ratio in the wind.

Oral contribution, published in ”Stellar Pulsation: Challenges for Theory and Observations”, Santa Fe, 2009 Available from arXiv:0908.4591

The enigma of B-type pulsators in the SMC

Sébastien Salmon^{1}, Josefina Montalbán^{1}, Andrea Miglio^{1}, Thierry Morel^{1}, Marc-Antoine Dupret^{1} and Arlette Noels^{1}

1Université de Liège, Belgium

Since the early nineties it is accepted that the excitation mechanism of B-type pulsators on the main sequence is due to the opacity peak in the iron-group elements at T ≈ 200, 000 K. The Fe content plays then a major role in the excitation of β Cep and SPB pulsations. While theoretical non-adiabatic computations predict no β Cep pulsators and only a small number of SPBs for low metallicity environments such as that of the Magellanic Clouds (MCs), recent variability surveys of B stars in the SMC have reported the detection of a significant number of SPB and β Cep candidates. Since SMC is the metal poorest (Z ≈ 0.001–0.004) of the MCs, it constitutes an interesting object for investigating the disagreement between theory and observations. We approach the problem by calling into question some of the hypotheses made in previous studies: given the different chemical evolution of the SMC compared with our local galactic environment, is it appropriate to describe the chemical composition of SMC B stars by scaling the solar mixture to lower Z? Is that composition uniform in space and time? In this paper we present the results of a stability analysis of B-type stellar models computed with a revised chemical composition and metallicity specific to the SMC.

Poster contribution, published in ”Stellar Pulsation: Challenges for Theory and Observation” (31 May – 5 June, Santa Fe, New Mexico, USA) Available from arXiv:0908.4185

Young open clusters in the Milky Way and Small Magellanic Cloud

Christophe Martayan^{1,2}

1European Organisation for Astronomical Research in the Southern Hemisphere, Alonso de Cordova 3107, Vitacura, Casilla 19001, Santiago 19, Chile 2GEPI, Observatoire de Paris, CNRS, Université Paris Diderot, 5 place Jules Janssen, 92195 Meudon Cedex, France

NGC 6611, Trumpler 14, Trumpler 15, Trumpler 16, Collinder 232 are very young open clusters located in star-formation regions of the Eagle Nebula or the Carina in the MW, and NGC 346 in the SMC. With different instrumentations and techniques, it was possible to detect and classify new Herbig Ae/Be stars, classical Be stars and to provide new tests / comparisons about the Be stars appearance models. Special stars (He-strong) of these star-formation regions are also presented.

Poster contribution, published in IAU Symposium 266 Available from arXiv:0909.3029

Evolution and appearance of Be stars in SMC clusters

C. Martayan^{1,2}, D. Baade^{3}, Y. Frémat^{4} and J. Zorec^{5}

1European Organisation for Astronomical Research in the Southern Hemisphere, Alonso de Cordova 3107, Vitacura, Casilla 19001, Santiago 19, Chile 2GEPI, Observatoire de Paris, CNRS, Université Paris Diderot, 5 place Jules Janssen, 92195 Meudon Cedex, France 3European Organisation for Astronomical Research in the Southern Hemisphere, Karl-Schwarzschild-Str. 2, 85748 Garching b. München, Germany 4Royal Observatory of Belgium, 3 avenue circulaire, 1180 Brussels, Belgium 5Institut d’Astrophysique de Paris, UMR7095, CNRS, Université Marie & Pierre Curie, 98bis Boulevard Arago 75014 Paris, France

Star clusters are privileged laboratories for studying the evolution of massive stars (OB stars). One particularly interesting question concerns the phases, during which the classical Be stars occur, which unlike HAe/Be stars, are not pre-main sequence objects, nor supergiants. Rather, they are extremely rapidly rotating B-type stars with a circumstellar decretion disk formed by episodic ejections of matter from the central star. To study the impact of mass, metallicity, and age on the Be phase, we observed SMC open clusters with two different techniques: 1) with the ESO-WFI in its slitless mode, which allowed us to find the brighter Be and other emission-line stars in 84 SMC open clusters 2) with the VLT-FLAMES multi-fiber spectrograph in order to determine accurately the evolutionary phases of Be stars in the Be-star rich SMC open cluster NGC 330. Based on a comparison to the Milky Way, a model of Be stellar evolution / appearance as a function of metallicity and mass / spectral type is developed, involving the fractional critical rotation rate as a key parameter.

Poster contribution, published in IAU Symposium 266 Available from arXiv:0909.3915

Job Advert

post-doctoral position

Background

VISTA is a 4-m class wide field survey telescope for the southern hemisphere, equipped with a large near infrared camera. VISTA is going through commissioning and science verification right now.

During the first few years of operation VISTA will exclusively be used to carry out previously selected “public surveys” that aim at addressing fundamental questions in astrophysics today. One of these is the VISTA Magellanic Cloud (VMC) survey, which will observe the entire Magellanic Cloud system (180 square degrees) in Y JK (for details see http://star.herts.ac.uk/^{∼}mcioni/vmc/). The main aims are to study the spatially resolved Star Formation History (SFH) and the 3D structure of and interaction between the MC subsystems.

At the Royal Observatory there is a vacancy for a postdoctoral researcher to work on the analysis of the tens of thousands late-type stars (AGB stars) that will be detected with VMC. Combining the VMC data with other photometric catalogs (optical, other NIR, Spitzer, Akari, Herschel) will allow their Spectral Energy Distributions to be constructed and then modelled in a quantitative way. In such a way a detailed study of the mass-loss process and AGB evolution as a function of environment over the MC will be possible for the first time.

The K-band data of VMC will be obtained over a 6 month period. This will allow the detection of Long Period Variables, and opens the possibility of combining this data with data from the MACHO and OGLE microlensing surveys. Studying the period distribution of LPVs/AGB stars over the MC will shed important light on the star formation history over the MC system.

Profile

PhD in astrophysics at the time the position is taken. Knowledge or familiarity with one or more of the following areas: AGB stars and their evolution, (dust) radiative transfer modelling, Virtual Observatory techniques (e.g. crosscorrelation), data reduction of near-infrared (survey) data.

Employment

Starting date: January 1^{st}, 2010, or as soon as possible thereafter. The position is initially for two years, but an extension of two years is possible. The duty station is the Royal Observatory of Belgium, which is situated in the south of Brussels. Salary and other benefits are according to the salary scales of the Federal Government.

Applications

The application, including a CV, publication list, description of research interest, and the names and contact details of two professional references that may be contacted, should be sent electronically to Dr. Martin Groenewegen (marting@oma.be). The position will remain open until filled, but applications received by November 15^{th} will receive first consideration.

See also http://homepage.oma.be/marting/Aktie1 vacature.pdf